a=1+b
b=5-2a
solve

Answer by checkley71(8403)

a=1-b & b=5-2a or b=5-2(1-b) or b=5-2+2b or b=-3 then a=1-(-3) or a=4
proof 4=1-(-3) or 4=4 & -3=5-2*4 or -3=-3